prove that A is a square matrix, then AAT and A + AT are symmetric


1
Expert's answer
2021-05-31T18:29:11-0400

A square matrix whose transpose is equal to itself is called a symmetric matrix; that is, B is symmetric if "B^T=B,B" is a square matrix.


Transpose of a product "(CD)^T=D^TC^T"

The operation of taking the transpose is an involution (self-inverse): "(B^T)^T=B"


The transpose respects addition: "(C+D)^T=C^T+D^T"


Then


"(AA^T)^T=(A^T)^TA^T=AA^T"

"AA^T" is symmetric.




"(A+A^T)^T=A^T+(A^T)^T=A^T+A=A+A^T"

"A+A^T" is symmetric.